# Patch-Level Training for Large Language Models

### v2

@@ 2 Patch-Level Training
While formulating the patch-level model structure, our goal is to minimize the discrepancy between patch-level and token-level models, thereby ensuring that the knowledge gained during patch-level training can be smoothly transferred to the token-level model. Given the context length $T$ for token-level training, we set the context length for patch-level training as $KT$, which is then compressed to a patch sequence of length $T$ to maintain consistency with the subsequent token-level training. To avoid introducing unnecessary parameters during token-to-patch compression, we represent the patch embedding as the average of its associated token embeddings. Let $p_{i}$ be the $i$-th patch, $x_{iK+k}$ be the $k$-th token in the $i$-th patch, and $E$ be the embedding function. The patch embedding is:
@@
The patch-level model is trained through next patch prediction, i.e., predicting the $K$ tokens in the next patch. The simultaneous prediction of multiple tokens has been explored in speculative decoding, which typically employs multiple output heads and each head is responsible for predicting a distinct token (Cai et al., 2024; Lin et al., 2024). However, this approach would also entail additional parameters that may be unfavorable for the subsequent knowledge transfer. Instead, we maintain a single output head and make its prediction cover all tokens in the next patch. Specifically, we calculate the cross-entropy loss for all the subsequent $K$ tokens based on the same patch prediction $p_{\theta}(\cdot|p_{<i})$, resulting in the following loss function:
$\mathcal{L}_{patch}=-\sum_{i=1}^{T}\sum_{k=0}^{K-1}\log p_{\theta}(x_{iK+k}|p_{<i}).$ (3)
@@ 3.2 Main Results
We train a series of LLMs of varying sizes (370M-2.7B parameters) on the Pile dataset. To halve the training costs, we employ patch-level training with the settings of $K=4,\lambda=2/3$, and compare its performance with the conventional token-level training. For the Transformer-370M, we also explore other choices of $\lambda$ to evaluate its impact. Table 1 presents the performance comparison of the resulting models. Remarkably, our approach consumes only half of the compute and incurs almost no performance loss. It matches the baseline model in terms of perplexity and even demonstrates a consistent gain in zero-shot evaluations, raising the average accuracy by approximately 0.5%. The model performance is also influenced by the choice of $\lambda$. Within the range of values we set, a smaller $\lambda$ leads to better model performance but also entails larger training costs. A more detailed study on the hyperparameter $\lambda$ will be presented in Section 3.6.
@@
Our primary motivation for patch-level training is to enhance the model’s knowledge acquisition efficiency. Interestingly, experimental results show that this approach can sometimes lead to performance improvements, which is beyond our initial expectation. We initially thought that the extended context length in patch-level training contributes to the improvements. However, when we decreased the context length during patch-level training from $KT=8192$ to $T=2048$ for Transformer-370M ($\lambda=1/2$), the model performance only experiences a slight decline (PPL $\uparrow$ 0.06, zero-shot accuracy $\downarrow$ 0.2), yet still surpasses the baseline, implying that context length is not the primary factor. We hypothesize that two other factors might be responsible for these improvements: firstly, the patch-level initialization could potentially serve as a form of regularization; secondly, by compressing tokens into patches, the model might more effectively recognize and capture long-range dependencies due to the reduced token distance.
@@ 3.4 Scaling Properties
In Table 1, we notice a trend of perplexity related to the model size: the performance advantage of patch-level training appears to decrease as the model size increases. Table 3 describes this trend more precisely, indicating that the model with patch-level training experiences smaller performance gains from the increase in model size. On the other hand, Table 4 presents the perplexity changes when maintaining a constant model size and varying the size of the training data. As the data size increases, the performance of patch-level training improves at a faster rate compared to the baseline model.
Table 3: Perplexity scores when scaling the model size from 370M to 2.7B and training on the Pile dataset (360B tokens). $\downarrow$’ indicates the perplexity reduction compared to the previous model size.
@@
Transformer 10.94 9.18 ($\downarrow$1.76) 8.18 ($\downarrow$1.0) 7.11 ($\downarrow$1.07)
+ Patch (${\lambda=2/3}$) 10.68 9.10 ($\downarrow$1.58) 8.23 ($\downarrow$0.87) 7.24 ($\downarrow$0.99)
Table 4: Perplexity scores of Transformer-370M when scaling the size of training data from 45B to 360B. $\downarrow$’ indicates the perplexity reduction compared to the previous data size. The batch size is adjusted to maintain a consistent number of training steps.
@@
Transformer 12.50 11.71 ($\downarrow$0.79) 11.28 ($\downarrow$0.43) 10.94 ($\downarrow$0.34)
+ Patch (${\lambda=2/3}$) 12.84 11.80 ($\downarrow$1.04) 11.17 ($\downarrow$0.63) 10.68 ($\downarrow$0.49)
@@ 3.5 Effect of Patch Size $K$
We investigate the effect of patch size under the settings of 90B training data, 370M model parameters, a batch size of 512K, and $\lambda=1/2$. The results are shown in Figure 5. Across different patch sizes, the loss curves for patch sizes $K=2$ and $K=4$ are nearly indistinguishable, while further increasing the patch size to 8 or 16 results in a certain performance decline. Despite this, these models still exhibit significant performance improvements when compared to the model trained from scratch, which does not benefit from the initialization of patch-level training.
@@ 3.7 Effect of Architecture
In the current setup, the architecture of the patch-level model is identical to that of the token-level model, facilitating smoother model transfer but also leading to some concerns. For instance, the token-to-patch transformation and the prediction of the next patch do not take into account the order of tokens within a patch. It has been suggested that patch-level models might benefit from adopting architectures that are better suited for modeling and predicting patches, where the additional parameters introduced in such architectures could simply be discarded when transferring the model to token-level training.
We evaluate this strategy under the settings of 90B training tokens, 370M model parameters, a batch size of 512K, and $\lambda=1/2$. Specifically, we incorporate a linear mapping layer at both the input and output sides of the model. On the input side, the conversion of token embeddings into patch embeddings is facilitated through a linear mapping $w_{in}\in\mathbb{R}^{Kd\times d}$. On the output side, a linear mapping $w_{out}\in\mathbb{R}^{d\times Kd}$ is employed to transform patch-level representations back into token-level representations, followed by a softmax layer to obtain the probability distribution of each token. The effects of these two modules are detailed in Table 5.
Table 5: Impact of architecture modifications in patch-level models. ‘+ InputMap’ and ‘+ OutputMap’ denote the incorporation of linear mappings at model’s input and output, respectively. ‘Patch PPL’ and ‘Token PPL’ are perplexities of the patch-level model and the final token-level model, respectively.
Model Transformer + InputMap + OutputMap + Both
@@
Overall, while these modifications are effective in reducing patch-level loss, they do not translate into benefits for the subsequent token-level training. Particularly, when both modifications are incorporated, although the patch-level loss is optimized to its lowest, the performance of the subsequent token-level model significantly declines. This suggests that there is no direct correlation between patch-level loss and the final performance of the token-level model, so a large loss during patch-level training does not imply ineffective learning. Therefore, we opt to preserve the original Transformer architecture and employ the most straightforward method for patch-level training.
@@ 4 Related Work
Multi-Token Prediction. Our approach improves training efficiency by concurrently predicting all tokens in the next patch. Similar attempts of multi-token prediction have been made in the past to improve the inference efficiency, including non-autoregressive generation (Gu et al., 2018) and speculative decoding (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023). Non-autoregressive generation reduces the number of decoding iterations by generating all tokens at once, involving techniques such as knowledge distillation (Kim & Rush, 2016; Shao et al., 2022), training objectives (Shao et al., 2021; Ghazvininejad et al., 2020; Du et al., 2021; Shao & Feng, 2022; Ma et al., 2023), latent modeling (Kaiser et al., 2018; Ma et al., 2019; Shu et al., 2020), iterative decoding (Lee et al., 2018; Ghazvininejad et al., 2019; Gu et al., 2019), and expressive model architectures (Libovický & Helcl, 2018; Huang et al., 2022; Gui et al., 2023). Despite the reduction in decoding iterations, the computational demands (FLOPs) do not decrease and may even rise for certain structures, leading to increased training costs. Speculative decoding is a novel decoding paradigm for accelerating LLM inference. During each step of decoding, it drafts several future tokens efficiently and then verifies them in parallel. The model for generating draft tokens can either be a relatively small model (Leviathan et al., 2023; Chen et al., 2023) or a non-autoregressive model that generates multiple tokens in parallel (Cai et al., 2024; Lin et al., 2024; Fu et al., 2024; Li et al., 2024). Recently, Gloeckle et al. pointed out that besides accelerating the inference, multi-token prediction also serves as an auxiliary training task to enhance the training signal. The key difference between our approach and these works lies in our utilization of multi-token prediction for reducing sequence length during training, with the goal of improving training efficiency. Regarding the model structure, we avoid introducing extra parameters and employ a single head for multi-token prediction.

### v3

# Beyond Next Token Prediction: Patch-Level Training for Large Language Models

@@ 2 Patch-Level Training
While formulating the patch-level model structure, our goal is to minimize the discrepancy between patch-level and token-level models, thereby ensuring that the knowledge gained during patch-level training can be smoothly transferred to the token-level model. In practice, we observed that it is crucial for at least one end (input or output) of the patch-level model to remain consistent with the token-level model, which acts as an anchor to align their representations. The detailed ablation is presented in section 3.7.
Given the context length $T$ for token-level training, we set the context length for patch-level training as $KT$, which is then compressed to a patch sequence of length $T$ to maintain consistency with the subsequent token-level training. To avoid introducing unnecessary parameters during token-to-patch compression, we represent the patch embedding as the average of its associated token embeddings. Let $p_{i}$ be the $i$-th patch, $x_{iK+k}$ be the $k$-th token in the $i$-th patch, and $E$ be the embedding function. The patch embedding is:
@@
The patch-level model is trained through next patch prediction, i.e., predicting the $K$ tokens in the next patch. The simultaneous prediction of multiple tokens has been explored in speculative decoding, which typically employs multiple output heads and each head is responsible for predicting a distinct token (Cai et al., 2024; Lin et al., 2024). However, this approach would also entail additional parameters that may be unfavorable for the subsequent knowledge transfer. Instead, we maintain a single output head and make its prediction cover all tokens in the next patch. Specifically, we calculate the cross-entropy loss for all the subsequent $K$ tokens based on the same patch prediction $P(\cdot|p_{<i})$, resulting in the following loss function:
$\mathcal{L}_{patch}=-\sum_{i=1}^{T}\sum_{k=0}^{K-1}\log P(x_{iK+k}|p_{<i}).$ (3)
@@ 3.2 Main Results
We train a series of LLMs of varying sizes (370M-2.7B parameters) on the Pile dataset. We employ patch-level training with the settings of $K=4,\lambda=2/3$, which theoretically reduces the training costs to 0.5×0.5$\times$0.5 ×. Please refer to Appendix B for the actual speed measurement. For the Transformer-370M, we also explore other choices of $\lambda$ to evaluate its impact. Table 1 presents the performance comparison of the resulting models. Remarkably, our approach consumes only half of the compute and incurs almost no performance loss. It matches the baseline model in terms of perplexity and even demonstrates a consistent gain in zero-shot evaluations, raising the average accuracy by approximately 0.5%. The model performance is also influenced by the choice of $\lambda$. Within the range of values we set, a smaller $\lambda$ leads to better model performance but also entails larger training costs. A more detailed study on the hyperparameter $\lambda$ will be presented in Section 3.6.
@@
Our primary motivation for patch-level training is to enhance the model’s knowledge acquisition efficiency. Interestingly, experimental results show that this approach can sometimes lead to performance improvements, which is beyond our initial expectation. We initially thought that the extended context length in patch-level training contributes to the improvements. However, when we decreased the context length during patch-level training from $KT=8192$ to $T=2048$ for Transformer-370M ($\lambda=1/2$), the model performance only experiences a slight decline (PPL $\uparrow$ 0.06, zero-shot accuracy $\downarrow$ 0.2), yet still surpasses the baseline, implying that context length is not the primary factor. We hypothesize that two other factors might be responsible for these improvements: firstly, the patch-level initialization could potentially serve as a form of regularization; secondly, by compressing consecutive tokens into patches, the model might more effectively recognize and capture long-range dependencies due to the reduced token distance.
@@ 3.4 Scaling Properties
In Table 1, we notice a trend related to the model size: the performance advantage of patch-level training appears to decrease as the model size increases. Table 3 describes this trend more precisely, indicating that the model with patch-level training experiences smaller performance gains from the increase in model size. On the other hand, Table 4 presents the changes of cross-entropy loss when maintaining a constant model size and varying the size of the training data. As the data size increases, the performance of patch-level training improves at a faster rate compared to the baseline model.
Table 3: Test losses when scaling the model size from 370M to 2.7B and training on the Pile dataset (360B tokens). $\downarrow$’ indicates the loss reduction compared to the previous model size.
@@
Transformer 2.392 2.217 ($\downarrow$0.175) 2.102 ($\downarrow$0.115) 1.961 ($\downarrow$0.141)
+ Patch (${\lambda=2/3}$) 2.368 2.208 ($\downarrow$0.160) 2.108 ($\downarrow$0.100) 1.980 ($\downarrow$0.128)
Table 4: Test losses of Transformer-370M when scaling the size of training data from 45B to 360B. $\downarrow$’ indicates the loss reduction compared to the previous data size. The batch size is adjusted to maintain a consistent number of training steps.
@@
Transformer 2.526 2.460 ($\downarrow$0.066) 2.423 ($\downarrow$0.037) 2.392 ($\downarrow$0.031)
+ Patch (${\lambda=2/3}$) 2.553 2.468 ($\downarrow$0.085) 2.413 ($\downarrow$0.055) 2.368 ($\downarrow$0.045)
@@ 3.5 Effect of Patch Size $K$
We investigate the effect of patch size under the settings of 90B training tokens, 370M model parameters, a batch size of 512K, and $\lambda=1/2$. The results are shown in Figure 5. Across different patch sizes, the loss curves for patch sizes $K=2$ and $K=4$ are nearly indistinguishable, while further increasing the patch size to 8 or 16 results in a certain performance decline. Despite this, these models still exhibit significant performance improvements when compared to the model trained from scratch, which does not benefit from the initialization of patch-level training.
@@ 3.7 Effect of Architecture
In the current setup, the architecture of the patch-level model is identical to that of the token-level model, facilitating smoother model transfer but also leading to some concerns. For instance, the token-to-patch transformation and the prediction of the next patch do not take into account the order of tokens within a patch. One may conjecture that patch-level models might benefit from adopting architectures that are better suited for representing and predicting patches, where the additional parameters introduced in such architectures could simply be discarded in the next stage.
We evaluate this strategy under the settings of 90B training tokens, 370M model parameters, a batch size of 512K, and $\lambda=1/2$. Specifically, we incorporate a linear projection layer at both the input and output sides of the model. On the input side, the conversion of token embeddings into patch embeddings is facilitated through a linear projection $w_{in}\in\mathbb{R}^{Kd\times d}$. On the output side, a linear projection $w_{out}\in\mathbb{R}^{d\times Kd}$ is employed to transform patch-level representations back into token-level representations, followed by $K$ softmax layers to obtain the probability distribution of each token. The effects of these two modules are detailed in Table 5.
Table 5: Impact of architecture modifications in patch-level models. ‘+ InputProj’ and ‘+ OutputProj’ denote the incorporation of linear projections at model’s input and output, respectively. ‘Patch PPL’ and ‘Token PPL’ are perplexities of the patch-level and token-level model, respectively.
Model Transformer + InputProj + OutputProj + Both
@@
Overall, while these modifications are effective in reducing the patch-level loss, they do not translate into benefits for the subsequent token-level training. Particularly, when linear projections are applied at both the model input and output, the performance of the subsequent token-level model significantly declines. It suggests that it is crucial for at least one end (input or output) of the patch-level model to remain consistent with the token-level model, which acts as an anchor to align their representations. It also shows that there is no direct correlation between patch-level loss and the final performance of the token-level model, so a large loss during patch-level training does not imply ineffective learning. Therefore, we opt to preserve the original Transformer architecture for patch-level training.
@@ 4 Related Work
Multi-Token Prediction. Our approach improves training efficiency by concurrently predicting all tokens in the next patch. Similar attempts of multi-token prediction have been made in the past to improve the inference efficiency, including non-autoregressive generation (Gu et al., 2018) and speculative decoding (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023). Non-autoregressive generation reduces the number of decoding iterations by generating all tokens at once, involving techniques such as knowledge distillation (Kim & Rush, 2016; Shao et al., 2022), training objectives (Shao et al., 2019; 2020; 2021; Ghazvininejad et al., 2020; Du et al., 2021; Shao & Feng, 2022; Ma et al., 2023; Shao et al., 2023), latent modeling (Kaiser et al., 2018; Ma et al., 2019; Shu et al., 2020), iterative decoding (Lee et al., 2018; Ghazvininejad et al., 2019; Gu et al., 2019), and expressive model architectures (Libovický & Helcl, 2018; Huang et al., 2022; Gui et al., 2023). Despite the reduction in decoding iterations, the computational demands (FLOPs) do not decrease and may even rise for certain structures, leading to increased training costs. Speculative decoding is a novel decoding paradigm for accelerating LLM inference. During each step of decoding, it drafts several future tokens efficiently and then verifies them in parallel. The model for generating draft tokens can either be a relatively small model (Leviathan et al., 2023; Chen et al., 2023) or a non-autoregressive model that generates multiple tokens in parallel (Cai et al., 2024; Lin et al., 2024; Fu et al., 2024; Li et al., 2024). Recently, Gloeckle et al. pointed out that besides accelerating the inference, multi-token prediction also serves as an auxiliary training task to enhance the training signal. The key difference between our approach and these works lies in our utilization of multi-token prediction for reducing sequence length during training, with the goal of improving training efficiency. Regarding the model structure, we avoid introducing extra parameters and employ a single head for multi-token prediction.
@@ 5 Conclusion
Appendix A Pseudocode
1# Model input
2num_patches = seq_length // self.patch_size
3inputs_embeds = inputs_embeds.view(batch_size, num_patches, self.patch_size, -1).mean(2)
4position_ids = position_ids[:, :num_patches]
5
6...
7
8# Model output
9shift_logits = logits[..., :-1, :].reshape(-1, self.config.vocab_size)
10shift_labels = labels[..., self.patch_size:].reshape(-1, self.patch_size)
11loss = 0
12log_probs = F.log_softmax(shift_logits, dim=1)
13for i in range(self.patch_size):
14 loss = loss + F.nll_loss(log_probs, shift_labels[:, i])
15loss = loss / self.patch_size
Appendix B Speed Measurement
The practical speed-up achieved through patch-level training falls short of the theoretical $K\times$ improvement, primarily due to overheads of data loading and gradient synchronization. Table 6 gives the actual running speed of patch-level training in comparison with the token-level baseline setting (patch_size=1, block_size=2048, per_device_train_batch_size=4, accumulation_steps=8), measured on 8 NVIDIA A100 GPUs. Increasing the patch size to 4 and the block length to 8192 results in a speed-up of approximately 3.8$\times$, albeit with some efficiency loss due to data loading. In practice, the gradient accumulation steps should be reduced to 2 to keep the total batch size constant, which brings the speed-up down to around 3.5$\times$ due to more frequent gradient synchronizations. Consequently, the actual runtime is approximately $\frac{2}{3}\times\frac{1}{3.5}+\frac{1}{3}\approx 0.523\times$, which is slightly larger than the theoretical cost of 0.5×0.5$\times$0.5 ×.
Table 6: Speed comparison between token-level and patch-level training.
Settings Tokens / sec Speed-up
Baseline 247.1K 1.0$\times$
patch_size=4, block_size=8192 933.5K 3.78$\times$
patch_size=4, block_size=8192, accumulation_steps=2 864.2K 3.50$\times$
Appendix C Mixup Training
The core idea of this work is to aggregate multiple tokens into a unit of higher information density, which we refer to as a ‘patch’. Our current strategy compresses every consecutive $K$ tokens into a single patch. Here we explore an alternative approach inspired by mixup data augmentation (Zhang et al., 2018). This method involves randomly selecting $K$ training samples and mixing their tokens position-wise to form a patch sequence. Similarly, we train the model with the next patch prediction objective.
We conduct experiments under the settings of 90B training tokens, 370M model parameters, a batch size of 512K, and $\lambda=1/2$. The results are shown in Figure 9. Mixup training offers some improvements over the model trained from scratch; however, its performance significantly lags behind that of patch-level training. The primary difference between these methods lies in the composition of the patches, suggesting that it is more beneficial to form patches from consecutive tokens. We hypothesize that by compressing consecutive tokens into patches, the model may more effectively recognize and capture long-range dependencies, thanks to the reduced distance between tokens.
Figure 9: Test losses of Transformer-370M w.r.t the number of processed tokens. Models are initialized by mixup training or patch-level training with patch size $K$.
Appendix D Effect of Tokenizer
Our approach enhances training efficiency by increasing the information density of text units and reducing their quantity. Another dimension that works similarly is the choice of tokenizer, as a tokenizer with a larger vocabulary typically represents the same data with fewer tokens. This raises an interesting question about how the efficacy of patch-level training is affected by the tokenizer. The LLaMA2 tokenizer, used in our initial experiments, has a vocabulary size of 32,000. Here, we extend our study to include the LLaMA3 tokenizer (Dubey et al., 2024), whose vocabulary size is 128,000, to observe whether our approach remains effective with tokenizers that have a higher compression rate.
We continue to use the Pile dataset for the training, which comprises approximately 300B tokens when encoded with the LLaMA3 tokenizer, corresponding to about 85% of the token count compared to the LLaMA2 tokenizer. Figure 10 depicts the loss curves for the models trained using patch-level training ($K=4,\lambda=2/3$) and from scratch at the token-level. Table 7 presents a performance comparison between the two approaches. Remarkably, patch-level training remains its effectiveness, achieving slightly better performance than token-level training while halving the training cost.
Figure 10: Negative log-likelihood (NLL) loss on test set w.r.t the number of processed tokens during the training of 370M-parameter Transformers. The LLaMA3 tokenizer is used.
Table 7: Performance comparison of Transformers trained on the Pile dataset. The LLaMA3 tokenizer is used.
Model Type Cost PPL MMLU HellaSwag PIQA WinoG ARC-E ARC-C Average
Transformer-370M 1.0$\times$ 15.48 22.9 41.2 67.3 54.1 45.6 25.8 42.8
+ Patch (${\lambda=2/3}$) 0.5$\times$ 15.32 23.1 41.1 68.1 53.2 45.9 26.2 42.9
The choice of patch size, we hypothesize, is also swayed by the tokenizer’s compression rate. When the token representations are less compact, a larger patch size might be employed to increase information density, and vice versa. We validate this hypothesis by comparing the effects of patch sizes $K=2$ and $K=4$ under the settings of 75B training tokens, 370M model parameters, a batch size of 512K, $\lambda=1/2$, and using the LLaMA3 tokenizer. The results are illustrated in Figure 11. It is observed that there is a slight performance gap between patch sizes $K=2$ and $K=4$ at this setting, whereas, in comparison, the loss curves for these patch sizes using the LLaMA2 tokenizer are almost indistinguishable in Figure 5. This confirms our hypothesis that the suitable patch size varies with the tokenizer. Nevertheless, we posit that a patch size of $K=4$ is still more advantageous under this setting, offering significant efficiency improvements at the cost of a minor performance trade-off.
Figure 11: Test losses of Transformer-370M w.r.t the number of processed tokens. Models are initialized by patch-level training with patch size $K$. The LLaMA3 tokenizer is used.